How many different ways can a manager and supervisor from a company branch in Manila be selected if there are 8 employees available?

Found 2 solutions by checkley79, stanbon:
Answer by checkley79(3341)

The first selection can be 1 of 8.
The second selection can be 1 of 7.
THUS:
8*7=56 different ways to select 2 positions from 8 candidates.

Answer by stanbon(75887)

How many different ways can a manager and supervisor from a company branch in Manila be selected if there are 8 employees available?
---
Because the 2 positions are not the same, order is implied.
-----
Ans: 8P2 = 8*7 = 56 ways
===========================
Cheers,
Stan H.
=============